OpenStudy (anonymous):

85/2m-92=23/2 (fraction eqation)

OpenStudy (anonymous):

all i need is the first part i can solve it from there

OpenStudy (aroub):

85/2m-92=32/2 ok first u multiply the whole equation by 2 to get rid of the 2 in the denominator so it becomes 85/m-92=23

OpenStudy (anonymous):

alright thanks

OpenStudy (aroub):

when i just saw ur msg i stoped and if u want to check ur answer i can give it to u so u could check if u have the same answer

OpenStudy (aroub):

m=17/23